# Task Diversity Shortens the In-Context Learning Plateau

### v2

@@ 6 Conclusion
Acknowledgement
We thank Wonjong Rhee for prompting this project as part of a course at Seoul National University. We are also grateful to the participants of the KAIST-SNU workshop for their discussions at the early stages of this work. Special thanks to Chanwoo Park for sharing his empirical insights. We thank Minkyu Kim for providing feedback about the figures. We also thank Hyunsik Chae, Junsu Kim, and Jaesung Park for reviewing the manuscript and providing valuable feedback.
Impact statement
This paper advances the understanding of training dynamics in multi-task training, contributing to the broader field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Experimental details
We focus on transformer (Vaswani et al., 2017), Mamba (Gu & Dao, 2024), and Hyena (Poli et al., 2023). To handle model architectures that process inputs and outputs as vectors in an embedding space, we first pad each $f(x_{i})$ with $(d-1)$ zeros to match the dimension of $x_{i}$s. We then add a learnable linear layer to map these vectors into the embedding space. A second learnable linear layer maps the model’s output to a scalar. Each model’s hyperparameters are dictated as follows. For transformer, we use GPT2 (Radford et al., 2019), with 12 decoder layers of embedding size 256 with 8 heads and relative positional encoding. Hyena follows the exact configuration of transformer’s setup for the relevant hyperparameters. On the other hand, Mamba employs 24 layers, doubling the number of layers to match the model parameter size as is standard for Mamba. All models contain input and output projection heads that map input to the hidden state and hidden state to output. To train a model from scratch, we randomly initialize the model parameters.
@@
The following pseudo algorithm summarizes the batch generation process of (multi)-task ICL training.
@@
Appendix B Language ICL tasks
B.1 Natural Language ICL tasks
@@
WordSelection$(d)$.
We devise WordSelection$(d)$ by generalizing the WordSelection Task (Fu et al., 2024). The goal of this task is to select a single word from given $d$ words. The model’s goal is to learn in-context which word to select.
@@
[…]
@@
Prior work Fu et al. (2024) reported the difficulty of learning WordSelection$(4)$, by showing that it cannot learn within about $78$k iterations. Indeed, as shown in (Figure 7, Right), single-task training of WordSelection$(4)$ exhibits a plateau for over $180$k iterations. However, mixing WordSelection$(4)$ and WordSelection$(2)$ significantly shortens the plateau to $3.2$k iterations, accelerating the learning. The same phenomenon is observed when WordSelection$(4)$ and WordLength are trained simultaneously. These results reveal that our claim generally holds in NLP tasks.
We also conduct the checkpoint experiment, where we use checkpoint models on WordSelection$(2)$ and WordLength as baseline models and train them respectively on WordSelection$(4)$. As shown in Figure 8, using checkpoint models shortens the plateau, demonstrating that NLP tasks also share a certain common structure.
B.2 Formal Language ICL task: Regbench
@@
Appendix C Checkpoint experiment
C.1 Experimental details
@@
Boolean Retrieval Linear Regression Quadratic Regression Sparse Linear Regression Gaussian Retrieval
Iteration number $105$k $19$k $12$k $7$k $50$k
@@
C.2 Experiment on Retrieval task
@@
Retrieval. Sample $1024$ 5-tuples of one key and four values: $\{(k_{i},v_{i}^{1},v_{i}^{2},v_{i}^{3},v_{i}^{4})\}_{i=1}^{1024}$. The $k_{i}$s and $v_{i}^{j}$s are independently sampled from $\mathcal{D}_{\mathcal{K}}$ and $\mathcal{D}_{\mathcal{V}}$, respectively. To generate prompts, we uniformly sample $(n-1)$ 5-tuples without replacement and uniformly choose one $v_{i}$ per 5-tuple, resulting in $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{n-1}$. Next, we sample $p\sim\mathrm{Unif\,}(\{1,\dots,n-1\})$ and set $\mathbf{q}=\mathbf{k}_{p}$. Finally, given $P_{n}=(\mathbf{k}_{1},\mathbf{v}_{1},\dots,\mathbf{k}_{n-1},\mathbf{v}_{n-1},\mathbf{q})$, the task is to predict $\mathbf{v}_{p}$. We consider two Retrieval tasks: Gaussian Retrieval and Boolean Retrieval, with $(\mathcal{D}_{\mathcal{K}},\mathcal{D}_{\mathcal{V}})=(\mathcal{N}(0,I_{d}),\mathcal{N}(0,1))$ and $(\mathcal{D}_{\mathcal{K}},\mathcal{D}_{\mathcal{V}})=(\mathrm{Unif\,}\{\pm 1\}^{d},\mathrm{Unif\,}\{\pm 1\})$, respectively.
@@
Appendix D Feature learning experiment
@@
Appendix E No-context learning
E.1 Optimal no-context functions
@@
E.2 Experimental details and additional figures
@@
for boolean ICL tasks. As shown by the following figures, these measurements are close to 00, which indicates that the model’s output corresponds to task-wise optimal no-context function. We follow Appendix A for other configurations. Moreover, to prevent the optimal no-context be merely a zero function, we use $\mu\neq 0$ to parameterize the in-context function classes.
Figure 13: No-context learning regime. During the plateau, the model output matches the task-wise optimal no-context function. (Top) Sparse Parity(2). (Middle) Quadratic Regression ($\mu=-1$). (Bottom) Quadratic Regression ($\mu=-1$) + Sparse Parity(2).

### v3

@@ 6 Conclusion
7 Acknowledgements
This research was supported by a grant from KRAFTON AI. We thank Wonjong Rhee for initiating this project as part of a course at Seoul National University. We are grateful to Junsu Kim, Jaesung Park, and Hyunsik Chae for their valuable feedback, and to the participants of the KAIST–SNU workshop for their insightful discussions at the early stages of this work. We also thank Chanwoo Park for sharing his empirical insights, and Minkyu Kim for providing feedback on the figures.
@@ Appendix A Experimental details
We focus on transformer (Vaswani et al., 2017), Mamba (Gu & Dao, 2024), and Hyena (Poli et al., 2023). To handle model architectures that process inputs and outputs as vectors in an embedding space, we first pad each $f(x_{i})$ with (d1)(d-1) zeros to match the dimension of $x_{i}$s. We then add a learnable linear layer to map these vectors into the embedding space. A second learnable linear layer maps the model’s output to a scalar. Each model’s hyperparameters are dictated as follows. For transformer, we use GPT2 (Radford et al., 2019), with 12 decoder layers of embedding size 256 with 8 heads and relative positional encoding. Hyena follows the exact configuration of transformer’s setup for the relevant hyperparameters. On the other hand, Mamba employs 24 layers, doubling the number of layers to match the model parameter size, as is standard for Mamba. All models contain input and output projection heads that map input to the hidden state and the hidden state to output. To train a model from scratch, we randomly initialize the model parameters.
@@
The following pseudo-algorithm summarizes the batch generation process of (multi)-task ICL training.
@@
Appendix B Training loss dynamics: One plateau, then rapid drop
In the ICL setups we consider, we observe that (i) there is a single plateau phase, and (ii) learning proceeds very rapidly once the model escapes this plateau. We provide several representative examples of training dynamics that support this observation below. For the cases of Sparse Parity(2) and Sparse Parity(3), training was performed jointly on a combination of three tasks: Sparse Parity(2), Sparse Parity(3), and Sparse Linear Regression.
Figure 8: Training loss curves showing a single plateau followed by rapid convergence.
Appendix C Detailed analysis of table 1
In Table 1, we provided comprehensive experimental evidence demonstrating that task diversity shortens the in-context learning (ICL) plateau, facilitating faster learning. To further robustly support our claims, we repeated each experiment three times with varying random seeds. The figure below plots the mean values along with standard deviations. It clearly shows a strong tendency: as the number of tasks increases, both the time to escape the plateau (blue) and the time to complete learning (orange) decrease significantly.
Figure 9: Sparse Parity(2)
Figure 10: Sparse Parity(3)
Figure 11: Linear Regression
Figure 12: Quadratic Regression
Figure 13: Leaky ReLu Regression
Figure 14: Sparse Linear Regression
Appendix D Language ICL tasks
D.1 Natural Language ICL tasks
@@
WordSelection(d)(d).
We devise WordSelection(d)(d) by generalizing the WordSelection Task (Fu et al., 2024). The goal of this task is to select a single word from given $d$ words. The model’s goal is to learn in-context which word to select.
@@
We also devise WordLength Task. Let $\text{len}(w)$ be the length of word $w$ and $\%$ be the remainder operator. Given four words $w_{1},\dots w_{4}$, the goal of this task is to learn in-context either $\sum_{i=1}^{4}\text{len}(w_{i})$ or $\sum_{i=1}^{4}\text{len}(w_{i})\%10$ from given examples.
@@
Prior work Fu et al. (2024) reported the difficulty of learning WordSelection(4)(4), by showing that it cannot learn within about 7878k iterations. Indeed, as shown in (Figure 7, Right), single-task training of WordSelection(4)(4) exhibits a plateau for over 180180k iterations. However, mixing WordSelection(4)(4) and WordSelection(2)(2) significantly shortens the plateau to 3.23.2k iterations, accelerating the learning. The same phenomenon is observed when WordSelection(4)(4) and WordLength are trained simultaneously. These results reveal that our claim generally holds in NLP tasks.
We also conduct the checkpoint experiment, where we use checkpoint models on WordSelection(2)(2) and WordLength as baseline models and train them respectively on WordSelection(4)(4). As shown in Figure 15, using checkpoint models shortens the plateau, demonstrating that NLP tasks also share a certain common structure.
D.2 Formal Language ICL task: Regbench
@@
Appendix E Checkpoint experiment
E.1 Experimental details
@@
Boolean
Retrieval
Linear
Regression
Quadratic
Regression
Sparse Linear
Regression
Gaussian
Retrieval
Iteration number 105105k 1919k 1212k 77k 5050k
@@
E.2 Experiment on Retrieval task
@@
Retrieval. Sample $1024$ 5-tuples of one key and four values: $\{(k_{i},v_{i}^{1},v_{i}^{2},v_{i}^{3},v_{i}^{4})\}_{i=1}^{1024}$. The $k_{i}$s and $v_{i}^{j}$s are independently sampled from $\mathcal{D}_{\mathcal{K}}$ and $\mathcal{D}_{\mathcal{V}}$, respectively. To generate prompts, we uniformly sample (n1)(n-1) 5-tuples without replacement and uniformly choose one $v_{i}$ per 5-tuple, resulting in $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{n-1}$. Next, we sample $p\sim\mathrm{Unif\,}(\{1,\dots,n-1\})$ and set $\mathbf{q}=\mathbf{k}_{p}$. Finally, given $P_{n}=(\mathbf{k}_{1},\mathbf{v}_{1},\dots,\mathbf{k}_{n-1},\mathbf{v}_{n-1},\mathbf{q})$, the task is to predict $\mathbf{v}_{p}$. We consider two Retrieval tasks: Gaussian Retrieval and Boolean Retrieval, with $(\mathcal{D}_{\mathcal{K}},\mathcal{D}_{\mathcal{V}})=(\mathcal{N}(0,I_{d}),\mathcal{N}(0,1))$ and $(\mathcal{D}_{\mathcal{K}},\mathcal{D}_{\mathcal{V}})=(\mathrm{Unif\,}\{\pm 1\}^{d},\mathrm{Unif\,}\{\pm 1\})$, respectively.
@@
Appendix F Feature learning experiment
@@
Appendix G No-context learning
G.1 Optimal no-context functions
@@
G.2 Experimental details and additional figures
@@
for boolean ICL tasks. As shown by the following figures, these measurements are close to 0, which indicates that the model’s output corresponds to task-wise optimal no-context function. We follow Appendix A for other configurations. Moreover, to prevent the optimal no-context be merely a zero function, we use $\mu\neq 0$ to parameterize the in-context function classes.
(a) Sparse Parity(2)
(b) Quadratic Regression ($\mu=-1$)
(c) QR ($\mu=-1$) + SP(2)
Figure 20: No-context learning regime. During the plateau, the model output matches the task-wise optimal no-context function.